# Birdie: Advancing State Space Models with Reward-Driven Objectives and Curricula

### v3

@@ Abstract
Efficient state space models (SSMs), such as linear recurrent neural networks and linear attention variants, offer computational advantages over Transformers but struggle with tasks requiring long-range in-context retrieval-like text copying, associative recall, and question answering over long contexts. Previous efforts to address these challenges have focused on architectural modifications, often reintroducing computational inefficiencies. In this paper, we propose a novel training procedure, Birdie, that significantly enhances the in-context retrieval capabilities of SSMs without altering their architecture. Our approach combines bidirectional input processing with dynamic mixtures of specialized pre-training objectives, optimized via reinforcement learning. We introduce a new bidirectional SSM architecture that seamlessly transitions from bidirectional context processing to causal generation. Experimental evaluations demonstrate that Birdie markedly improves performance on retrieval-intensive tasks such as multi-number phone book lookup, long paragraph question-answering, and infilling. This narrows the performance gap with Transformers, while retaining computational efficiency. Our findings highlight the importance of training procedures in leveraging the fixed-state capacity of SSMs, offering a new direction to advance their capabilities. All code and pre-trained models are available at https://www.github.com/samblouir/birdie, with support for JAX and PyTorch.
@@ 3.1 Bidirectional processing
Bidirectional processing has shown advantages in generative Transformers using Prefix language Modeling and Span Corruption (Infilling objectives (Raffel et al., 2020; Tay et al., 2023c; Dong et al., 2019). There is also a rich history of encoder-decoder nonlinear RNNs that compress a prefix into a single vector before generating the suffix (Kalchbrenner and Blunsom, 2013; Cho et al., 2014; Sutskever et al., 2014). Bidirectionality may enable SSMs to better triage state capacity, which is crucial for retrieval-intensive tasks. Our results indicate that bidirectional SSMs outperform their unidirectional counterparts on several such tasks.
To address these challenges, we adapt the bidirectional ’prefix-LM’ architecture proposed in T5 and UniLM to SSMs (Raffel et al., 2020; Dong et al., 2019). This adaptation allows us to match a standard causal configuration in both parameter count and compute. Since our models do not use Attention, we divide the recurrent state into forward and reverse components. The forward components are processed without modification, enabling our bidirectional layers to transmit information from the prefix to the suffix via the forward state dimensions. This contrasts with the bidirectional Encoder layers in Encoder-Decoder models, which are constrained to operate only within the prefix. The reverse components are modified in the suffix region to maintain causality; specifically, we zero out the forget gate ($A_{t}$) dimensions. This approach prevents the state from propagating information backward in causal areas. A mathematical description follows [^2] Efficient implementations are provided in our codebase: https://www.github.com/samblouir/birdie, and an additional example is included in Appendix Section A.4.
@@
We note that concurrent work, Arora et al. (2024b), also considers applying bidirectional processing to the prefix of linear attention models, similar to prior encoder-decoder works  Kalchbrenner and Blunsom (2013); Cho et al. (2014); Sutskever et al. (2014). These models use separate components and stages for processing the prefix and generating the target sequences. In contrast, our approach adapts a decoder-only, prefix-LM setup similar to that described in  Raffel et al. (2020); Tay et al. (2023c); Dong et al. (2019) but adapted to recurrent models. Additionally, Arora et al. (2024b) considers a fixed pre-training objective that mixes Prefix language Modeling with Masked Language Modeling, as in work such as UniLM and AlphaCode (Dong et al., 2019; Li et al., 2022). Our experimental results with the UL2 objective in Section 4 suggest fixed objectives such as this may be insufficient to improve SSM performance on retrieval-intensive tasks.
@@ 3.3 Optimal Mixtures with Objective Sampling via Reinforcement Learning
We visualize loss, greedy-decoding accuracy, and sampling probabilities for training objective categories in Figure 5 in Appendix A. We observe trends, such as the observation that training on the Autoencoding objective appears to boost both Copying and Deshuffling objectives to the extent that their sampling can be nearly shut-off. Other behaviors emerge, such as the selective copying ability continuing to form once the model sees sufficient amounts of these samples.
We combine this approach with the new objectives described in Section 3.2 and the bidirectional processing described above in Section 3.1 and Appendix Section A.3, and refer to the resulting method as Birdie. We observe that Birdie consistently improves SSM performance on a variety of downstream tasks, as related in Section 4.
@@ 3.4 Gated SSM baseline
In our basic Gated SSM above, we fuse the SSM and MLP blocks as is done in Gated State Spaces and Mamba (Mehta et al., 2023; Gu and Dao, 2023). We find this simple baseline performs comparably with state-of-the-art SSMs, such as Hawk, on max-likelihood tasks, but does not perform as well when asked to retrieve multiple phone numbers simultaneously, or when generating responses to questions about Wikipedia excerpts in SQuAD-V2.
@@ A.1.1 Objective Classes
  • Autoencoding (Deshuffling is included)
@@ A.4 Bidirectional Python Implementation Example:
We provide an efficient implementation of our approach on our Github page in PyTorch and JAX at https://www.github.com/samblouir/birdie. To further clarify our approach, we provide the following pseudocode in Python:
@@ A.6 UL2
UL2 is a mixture-of-denoisers found via ablations on Transformer models Tay et al. (2023b).
It uses the following mixture of infilling and prefix-language modeling objectives:
A.7 Denoising Objectives
We summarize the denoising objectives used in UL2.
@@ A.7 Denoising Objectives
”S” Denoisers Prefix language Modeling, sequence split 50% [^5] We note a minor discrepency between Table 1 in UL2’s paper uses an average of 75% of the input sample for the context. We use 50%, since this is what the provided code in UL2 does.
@@
This table organizes the denoisers by type, indicating specific parameters used in each denoising task. ”S” denoisers use a Prefix language Modeling approach where input samples are randomly split, typically at a 50% rate, for context generation.
UL2 prepends data samples using a ”paradigm” token that mark which objective a sample is noised with. We take a similar approach in Birdie, but, rather than always prepend the paradigm token to the context region, we also append the token 50% of the time.
UL2’s mixtures out-performed training solely on Next Token Prediction strongly on the majority of downstream tasks considered by its authors. On tasks where UL2 did not out-perform Next Token prediction, it was usually only by a small amount. PaLM-2’s procedure was not disclosed.
With Birdie, we aim to address the following limitations with UL2:
@@
    UL2’s training objective selection and ratios was ablated for Transformers, not recurrent models.
@@
    UL2’s objective selections do not appear to induce robust copying or retrieval abilities in SSMs. Our Gated SSM pre-trained using UL2 was not able to retrieve phone numbers any better than the Next Token Prediction model. However, the UL2 model was able to retrieve answers better than Next Token Prediction model in the SQuAD V2 Question Answering task.
@@
    UL2 uses fixed sampling ratios of objectives throughout the training process. The SSMs struggle on Span Corruption relative to the Transformers, they may need a greater sampling ratio for similar performance. Adjusting these ratios to be optimal for each model likely require costly ablations. Instead, we avoid ablations on our end by approximating optimal ratios using Birdie.

### v5

@@ Abstract
Efficient state space models (SSMs), such as linear recurrent neural networks and linear attention variants, offer computational advantages over Transformers but struggle with tasks requiring long-range in-context retrieval-like text copying, associative recall, and question answering over long contexts. Previous efforts to address these challenges have focused on architectural modifications, often reintroducing computational inefficiencies. In this paper, we propose a novel training procedure, Birdie, that significantly enhances the in-context retrieval capabilities of SSMs without altering their architecture. Our approach combines bidirectional input processing with dynamic mixtures of specialized pre-training objectives, optimized via reinforcement learning. We introduce a new bidirectional SSM architecture that seamlessly transitions from bidirectional context processing to causal generation. Experimental evaluations demonstrate that Birdie markedly improves performance on retrieval-intensive tasks such as multi-number phone book lookup, long paragraph question-answering, and infilling. This narrows the performance gap with Transformers, while retaining computational efficiency. Our findings highlight the importance of training procedures in leveraging the fixed-state capacity of SSMs, offering a new direction to advance their capabilities. All code and pre-trained models are available at https://www.github.com/samblouir/birdie.
@@ 3.1 Bidirectional Processing
Encoder-Decoder models encode a prefix sequence into a single vector (or set of hidden states, for cross-attention) to generate a suffix from (Kalchbrenner and Blunsom, 2013; Cho et al., 2014; Sutskever et al., 2014; Raffel et al., 2020). The prefix-LM architecture, introduced in T5 and UniLM (Raffel et al., 2020; Dong et al., 2019), simplifies this two-stage approach by instead only using self-attention layers with specific masking to allow for bidirectional processing in the prefix and causality in the suffix regions. We adapt this prefix-LM approach for recurrent models, which allows us to transfer information between the prefix and suffix regions for every layer, in contrast to the single-vector bottleneck with Encoder-Decoders.
This adaptation ensures our bidirectional SSMs maintain computational and parameter efficiency, enabling fair comparisons with unidirectional models.
To adapt prefix-LM to our models, we first split the recurrent state into forward and reverse components. The forward components are processed without modification, enabling our bidirectional layers to transmit information from the prefix to the suffix via the forward state dimensions. This contrasts with the bidirectional Encoder layers in Encoder-Decoder models, which are constrained to operate only within the prefix. The reverse components are modified in the suffix region to maintain causality; specifically, we zero out the forget gate ($A_{t}$) dimensions. This approach prevents the state from propagating information backward in causal areas. A mathematical description follows [^2] Efficient implementations are provided in our codebase: https://www.github.com/samblouir/birdie, and an additional example is included in Appendix Section A.4.
@@
Compared to causal SSMs, our bidirectional SSMs trained with Birdie show greatly enhanced capabilities on question-answering retrieval with information-dense Wikipedia articles in SQuAD V2, shown in Figure 3. As we allocate half of the state for the reverse direction, we conservatively only allow half of Birdie’s layers to be bidirectional. On the multi-phone number retrieval task, we only observe a small reduction In contexts where the information is Although Birdie - Causal allows for more information to flow forward
Performance is on-par with the Birdie - Causal model in the bandwidth […] Our findings demonstrate that these bidirectional SSMs surpass their unidirectional equivalents in multiple such tasks, underscoring the utility of bidirectional processing.
We note that concurrent work, Arora et al. (2024b) also applies bidirectional processing to the prefix area with linear attention models and applies independent weights for the prefix and suffix regions. Additionally, they consider a fixed pre-training objective that mixes Prefix language Modeling with Masked Language Modeling, as in work such as UniLM and AlphaCode (Dong et al., 2019; Li et al., 2022). Our experimental results with the UL2 objective in Section 4 suggest that objectives such as this may be insufficient to improve SSM performance on retrieval-intensive tasks.
@@ 3.3 Optimal Mixtures with Objective Sampling via Reinforcement Learning
We visualize loss, greedy-decoding accuracy, and sampling probabilities for training objective categories in Figure 5 in Appendix A. We observe several trends, such as that training on the Autoencoding objective appears to boost both the Copying and Deshuffling objectives to the extent that their sampling can be nearly shut-off. Other behaviors emerge, such as the selective copying ability continuing to form regardless of sampling rate after the model has seen sufficient amounts of these samples.
We integrate our automated approach with the new objectives introduced in Section 3.2 and the bidirectional processing from Section 3.1 and Appendix A.3, resulting in a method we call Birdie. As shown in Section 4, Birdie consistently enhances SSM performance across a variety of downstream tasks.
@@ 3.4 Gated SSM baseline
In our basic Gated SSM above, we fuse the SSM and MLP blocks as is done in Gated State Spaces and Mamba (Mehta et al., 2023; Gu and Dao, 2023). We find this simple baseline performs comparably with state-of-the-art SSMs, such as Hawk, on max-likelihood tasks, but does not perform as well when asked to retrieve multiple phone numbers simultaneously, or when generating responses to questions about Wikipedia excerpts in SQuAD V2.
@@ A.1.1 Objective Classes
  • Autoencoding with Deshuffling
@@ A.4 Bidirectional Python Implementation Example:
We provide an efficient implementation of our approach on our Github page in PyTorch and JAX at https://www.github.com/samblouir/birdie. To further clarify our approach, consider the following pseudocode using Python:
@@ A.6 UL2
”S” Denoisers Prefix language Modeling, sequence split 75%
@@
Paradigm Token Mean Span Width Masked Input %
”R” dDenoiser 3 tokens 15%
”X” dDenoiser 32 tokens 50%
”S” dDenoiser Prefix language Modeling, sequence split 75%
- Next Token Prediction (also referred to as Causal language Modeling)(
Table 6: Summary of Denoising Objectives used in a specialized variant of UL2 for decoder-only models, as reported in Garcia et al. (2023). The objectives are sampled at a ratio of 20% prefix language modeling, 60% next token prediction, and 20% infilling. Presumably, this translates to a 10% sampling probability of the ”R” dDenoiser and a 10% sampling probability for the ”S” denoiser. Note: a denoiser paradigm token was not provided for the next token prediction task, so. Thus, we do not include one here. More details are provided in subsection A.6
UL2, introduced by Tay et al. Tay et al. (2023b), is a mixture-of-denoisers (MoD) pre-training strategy originally developed and ablated for Transformer-based language models. Rather than relying solely on a causal (left-to-right) language modeling objective, UL2 interleaves multiple forms of denoising tasks, such as infilling and prefix language modeling, to produce more robust representations.
In its original formulation, UL2 labels each sequence with a “paradigm token” (e.g., “R” or “X”) to indicate which denoising objective was used. Table 5 summarizes the full set of infilling and prefix modeling objectives in UL2, including their mean span widths and the percentage of tokens masked. These paradigm tokens were ablated off by continuing pre-training without them, for a significant number of steps.  Garcia et al. (2023) published an updated and specialized variant for decoder-only models from the authors of UL2. This variant augments UL2 with a pure next token prediction objective (standard causal language modeling) and simplifies the amount of infilling configurations. Table 6 lists these modified denoisers, alongside their sampling ratios. In this variant, “R” denotes a short-span infilling objective, “X” denotes a more extreme infilling setup with higher mask percentages or larger spans, and “S” indicates a prefix language modeling task.
Empirical results in both the original and specialized versions show that mixing multiple denoising tasks can outperform training solely on next token prediction for many downstream tasks. In cases where UL2 does not surpass standard causal modeling, the authors noted performance differences tend to be small Tay et al. (2023b).
A.7 Denoising Objectives and Birdie Motivation
Tables 5 and 6 provide an overview of the denoising strategies used in UL2. The objective categories, “R,” “X,” “S,” (as well as next token prediction), are used to group specific noising objective and configuration classes that are applied to input samples. For instance, “R” often denotes standard span infilling at normal masking probabilities, while “X” can span larger regions of text and/or mask up to 50% of tokens. In the case of “S” (prefix language modeling), the sequence is partially visible at the beginning, requiring the model to continue predicting from a given prefix.
In our approach with Birdie, we follow a similar paradigm-token strategy, but introduce two key modifications:
@@ A.7 Denoising Objectives and Birdie Motivation
    Appending Paradigm Tokens Instead of always prepending the paradigm token, Birdie appends it 50% of the time. This design choice is intended to familiarize the model with varied token placements and reduce reliance on position-based cues.
@@
    Dynamic Mixture Ratios. While UL2 maintains fixed ratios of denoising tasks throughout training, Birdie adaptively updates the ratio of objectives (e.g., infilling vs. prefix vs. causal language modeling) based on model feedback signals. This avoids exhaustive ablation studies and aims to optimize performance for different architectures and tasks.
These modifications address several limitations we observed when applying UL2 to state-space models (SSMs):
  1. 1.
    Transformer-Centric Ablations. UL2’s reported ablations focus primarily on Transformers. Our experiments indicate that SSM-based models (e.g., Gated SSMs) respond differently to certain denoising objectives, necessitating an approach tailored to SSMs.
  2. 2.
    Weak Retrieval/Copy Abilities. Although UL2 excels on many downstream tasks, we found that SSMs trained with UL2 did not improve on exact-retrieval tasks (e.g., phone number copying) over a standard next-token predictor. This suggests the need for adaptive sampling of denoisers that more strongly emphasize copying skills for SSM architectures.
@@
    Fixed Objective Ratios. UL2 uses a constant mixture of objectives across the entire training process. Empirically, we observed that certain objectives (e.g., infilling) may require a higher sampling frequency for SSMs, whereas Transformers are able to perform this task relatively easy. Dynamic ratio adjustments in Birdie allow the model to re-balance objectives on the fly, rather than relying on predefined proportions.
Our approach is able to maintain the benefits of mixture-of-denoiser pre-training, while addressing our practical compute constraints and unique needs posed by SSMs.